OpenStudy (anonymous):

Can someone verify: x=(1/2)t , y=3t^3-1 =y=3/8t^3-1

OpenStudy (anonymous):

@Psymon

OpenStudy (psymon):

In another problem, but I'll look in a sec:3

OpenStudy (anonymous):

"y=3/8t^3-1" ? this is wrong

OpenStudy (anonymous):

my options are y = three divided by eight t3 - 1 y = 24x3 - 1 y = 6x3 - 1 y = 24x - 1 @Coolsector

OpenStudy (anonymous):

so from x= (1/2) * t you get t = 2x

OpenStudy (anonymous):

now plug it into y=3t^3-1 what do you get ?

OpenStudy (anonymous):

y=6x3-1?

OpenStudy (anonymous):

no: y= 3 * (2x)^3 - 1

OpenStudy (anonymous):

now (2x)^3 = 2^3 * x^3 ..

OpenStudy (anonymous):

but wouldn;t i multiply 2 and 3 to give my 6x^3?

OpenStudy (anonymous):

but power precede multiplication

OpenStudy (anonymous):

so 24?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

WAIT is it y=24x-1 or 24x^3-1?

OpenStudy (anonymous):

@Coolsector

OpenStudy (anonymous):

you can answer it.. look at (2x)^3